Q.A particle travels along a straight line acc to equation x=t(t-2).Find distance travelled between t=0sec and t=4sec.
Ans=2m

Grade:11

1 Answers

Eshan
5 years ago
Dear stuent,

v=\dfrac{dx}{dt}=2(t-1)
The particle turns at t=1s.

Hence displacent from t=0 to t=1s isx(t=1s)-x(t=0s)=-1m

Displacement from t=1s to t=4s isx(t=4s)-x(t=1s)=9m
Hence total distance travelled=10m.

Answer cannot be 2m. Please check.

Regards.